OpenStudy (anonymous):

If a line crosses the y-axis at (0, -4) and has a slope of -2, what is the equation of the line?

OpenStudy (anonymous):

2x − y = -4 2x + y = -4 2x + y = 4 2x − y = 4

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -2 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)